OpenStudy (anonymous):

Help DBQ: Find solution of the given initial value problem in explicit form: y'= (e^-x - e^x)/(3+4y), y(0)=1

OpenStudy (perl):

this equation can be separated

OpenStudy (perl):

hint, multiply both sides by 3 + 4y

OpenStudy (anonymous):

Yes and I got 3y+2y^2= -e^-x-e^x + c

OpenStudy (anonymous):

but I dont know where to go from there...

OpenStudy (anonymous):

yeah but then you have to take the integral of that, no?

OpenStudy (abb0t):

@Mimi_x3 can answer this for you.

OpenStudy (mimi_x3):

Wow wow ....

OpenStudy (perl):

ok i agree dan

OpenStudy (perl):

3y+2y^2= -e^-x-e^x + c , now plug in your initial condition

OpenStudy (anonymous):

y(0)=1?

OpenStudy (perl):

y(0) = 1 3*1 + 2*1^2 = -e^(-0) - e^(0) + C solve for c

OpenStudy (anonymous):

5/-2 = C?

OpenStudy (perl):

3 + 2 = -1 -1 + c 5 = -2 + c 5 + 2 = c

OpenStudy (anonymous):

lolz yeah 7.

OpenStudy (anonymous):

dont i have to do something with the quadratic formula?

OpenStudy (perl):

nope

OpenStudy (perl):

it will just get messy if you do

OpenStudy (perl):

hmm, it says give solution in 'explicit form'

OpenStudy (perl):

well this solution is explicit

OpenStudy (perl):

if your teacher specifically wants you to find a solution y(x)=f(x) , then yes you would have to use quadratic formula

OpenStudy (perl):

here is the solution y(x)